A photon of wavelength 4 angstrom an electron at rest and this photon is deflected at an angle of 1500 to its original direction. Determine the wavelength of this photon after the collision
1
Expert's answer
2021-08-31T17:06:16-0400

The wavelength of scattered photon is

"\lambda =\lambda _0+(\frac{h}{mv})(1-cos150^o)"

"=(4\times 10^{-10})+ \frac{(6.626\times 10^{-34})}{(9.11\times 10^{-31}(3\times 10^8)}(1-cos150^o)"

"=4.0452\times 10^{-10} m= 4.0452 Angstrom"